OpenStudy (anonymous):

6 10 % × 45 = _____ Select one of the options below as your answer: A. .027 B. .27 C. 2.7 D. 27

OpenStudy (anonymous):

6/10=0.6 0.6/100=0.006 0.006*45=0.27

OpenStudy (anonymous):

have more confidence in yourself

OpenStudy (anonymous):

Am I Right

OpenStudy (anonymous):

yes you are

OpenStudy (anonymous):

So it Woould Be B.27 Right

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

on a side note, please write your fractions as 1/2 or 2/3, or in this case 6/10

OpenStudy (anonymous):

and then surround it with a parenthesis